### v1

# FollowIR: Evaluating and Teaching Information Retrieval Models to Follow Instructions

Abstract
Modern Large Language Models (LLMs) are capable of following long and complex instructions that enable a diverse amount of user tasks. However, despite Information Retrieval (IR) models using LLMs as the backbone of their architectures, nearly all of them still only take queries as input, with no instructions. For the handful of recent models that do take instructions, it’s unclear how they use them. We introduce our dataset FollowIR, which contains a rigorous instruction evaluation benchmark as well as a training set for helping IR models learn to better follow real-world instructions. FollowIR builds off the long history of the TREC conferences: as TREC provides human annotators with instructions (also known as narratives) to determine document relevance, so should IR models be able to understand and decide relevance based on these detailed instructions. Our evaluation benchmark starts with three deeply judged TREC collections and alters the annotator instructions, re-annotating relevant documents. Through this process, we can measure how well IR models follow instructions, through a new pairwise evaluation framework. Our results indicate that existing retrieval models fail to correctly use instructions, using them for basic keywords and struggling to understand long-form information. However, we show that it is possible for IR models to learn to follow complex instructions: our new FollowIR-7B model has significant improvements (over 13%) after fine-tuning on our training set.
1 Introduction
Large language models (LLMs) have shown impressive results in understanding long complex instructions and have enabled a broad array of tasks (Brown et al., 2020; Achiam et al., 2023). The Information Retrieval (IR) community has adopted the use of LLMs as the main backbone of their neural retrieval architectures and has seen an explosion of progress in semantic search (Karpukhin et al., 2020; Khattab & Zaharia, 2020; Reimers & Gurevych, 2019). However, these existing systems rely on under-specified keywords and short sentences, despite the growing capabilities of LLMs. And even though the retrieval community is increasingly adopting larger IR models for their backbones (such as LLaMA 7B in Pradeep et al. (2023); Ma et al. (2023); Weller et al. (2023)), they are only used to estimate document relevance without any instructions. Moving past this ad-hoc search to search with instructions would enable experts to narrow in on complex information needs using only flexible natural language, adapting to new tasks on the fly. For example, imagine a researcher seeking to identify papers that must contain numerous qualities to be relevant (from a given venue, using a particular method, etc.) while also making sure to avoid conditions that would make the document not-relevant (on the wrong domain, from the wrong time period, etc.).
Recent work has started to move towards search with instructions, but is still understudied, with only a handful of papers on the topic (Muennighoff et al., 2024; Su et al., 2022; Asai et al., 2022). Although these works have shown great initiative in adapting retrieval models to take instructions, these instructions are typically short (less than 10 words), repetitive (only one instruction per dataset e.g. Su et al. (2022); Asai et al. (2022); Wang et al. (2023); Xiao et al. (2023)), and lack evaluation datasets that explicitly measure instruction following – instead using standard retrieval benchmarks.
To address this we introduce FollowIR, which contains a benchmark that explicitly measures the instruction following ability of retrieval models, while also providing training data that includes diverse and realistic instructions. Our key intuition is that many ad-hoc retrieval collections developed as part of TREC (Text REtrieval Conference) conferences provide instructions to human annotators who judge document relevance for a given query, as TREC recognizes that information needs often cannot be specified in only a simple query. These instructions are thorough and complex, including minute details about what makes a document relevant vs not-relevant. Thus if humans can use these TREC instructions to annotate document relevance, so should instruction-following retrieval models (example query and instruction pairs are shown in Figure 1 and Table 1).
We use three deeply-judged111i.e. that a large number of documents have been judged relevant or non-relevant, see Section 2.1 for more. TREC collections as the basis of our evaluation set: TREC Robust 2004 (Voorhees, 2005), TREC Common Core 2017 (Allan et al., 2017), and TREC News 2021 (Soboroff et al., 2020). These collections have been thoroughly annotated in order to evaluate recall in retrieval, frequently with hundreds (or even thousands) of documents judged as relevant or not-relevant. We take the instructions given to human annotators (called narratives) and alter them slightly, re-annotating the relevant documents. We then have paired instructions, which can be used to test how models respond to changing instructions – seeing if they change the set of relevant documents to correspond to the relevance definition in the new altered instructions.
Feature Description Example %
N/A: Query-only
Used for keyword and semantic similarity matching What impact has the Chunnel had on the British economy and/or the life style of the British?
Specifics Detailed specifics on what qualifies as relevant, potentially including or or and conditions Documents discussing the following issues are relevant: - projected and actual impact on the life styles of the British - Long term changes to economic policy and relations - major changes to other transportation systems linked with the Continent … 90.4
Irrelevant Info Not keyword relevant, but helpful for background … The Channel Tunnel, sometimes referred to informally as the Chunnel, is a 50.46-kilometre undersea railway tunnel, opened in 1994 … 34.6
Negation Specifics on what type of documents are not relevant and should be excluded … Documents discussing the following issues are not relevant: - expense and construction schedule - routine marketing ploys by other channel crossers (i.e., schedule changes, price drops, etc.) … 52.9
Table 1: How standard retrieval queries differ from instructions (or narratives): instructions contain more specific details about what is relevant, include less directly-relevant background information, and often have directives about what documents are not relevant, using negation. % indicates the percent of the time these features are present in the original TREC instructions used in FollowIR.
As there are no existing methods to compare pairwise queries in IR, we develop a new evaluation framework to do so, measuring rank-wise score changes (which we call $p$-MRR) of documents given a pair of different instructions with the same query. Results on FollowIR indicate that current models generally fail to follow instructions in retrieval unless they are 3B+ parameters or have not been trained for retrieval. Our analysis shows that these failures are due to two phenomena: the length of the instruction (as models are not used to long instructions) as well as using the instruction for keyword search rather than as an instruction that defines relevance.
To further progress in building retrieval models that can understand instructions, we build a training set of real-world human-used instructions and fine-tune a model on them (FollowIR-7B). Our results show marked improvement on FollowIR for both standard IR metrics and for $p$-MRR, indicating a starting point for future progress on instruction following.
Thus, we provide the following contributions: (1) a benchmark for evaluating instruction following in retrieval (FollowIR) consisting of human annotations on top of three already highly-judged corpora, (2) analysis of why current models fail to understand instructions, and (3) training data for teaching retrieval models to follow instructions along with a new open-sourced IR model, FollowIR-7B, that can handle long instructions in IR. [^2] Links to the code, data, and models are available at https://github.com/orionw/FollowIR
Figure 1: A visual depiction of the pairwise evaluation framework: models are evaluated on both the original query and instruction, as well as the altered instruction. If the model correctly understands the instructions, it will change which documents are relevant w.r.t. the alteration (right). Note that the real-world instructions (left) given to TREC annotators includes fine-grained details about what relevance is, as well as instructions containing negation (in bold).
2 Related Work
2.1 TREC Conferences
The United States National Institute of Science and Technology (NIST) created the TREC organization in 1993. Each year TREC sponsors many tracks, or shared tasks, on a given dataset. These tracks range from a variety of topics: anywhere from standard ad-hoc retrieval on news (Soboroff et al., 2018; Soboroff, 2021) to more complex domains such as legal retrieval (Baron et al., 2006; Oard et al., 2008) or retrieval-augmented generation/report-generation (Lawrie et al., 2024).
As part of this process, NIST sponsors annotations for these collections. Typically, this is done by pooling a set of results (runs) from a variety of retrieval models and then annotating them in rank order until funding runs out. To help facilitate annotation, track organizers provide a narrative (or instruction) for each query that will be given to the annotators – however, IR models are only ever given the query. As evaluating total recall would require annotating every document in the collection for every query (which is not feasible for collections with millions of documents), recall error is tested using post-hoc sampling and annotation. Although not every query and document pair can be evaluated, recall for queries is very high. We build off the rigorous evaluation done at TREC by using several of their collections to build FollowIR.
2.2 Instructions for LLMs
Instruction-following LLMs have been popularized by models such as GPT-3 (Brown et al., 2020) and have become a large area of interest for the natural language processing community (Touvron et al., 2023a; Jiang et al., 2023; Groeneveld et al., 2024; Black et al., 2022). There has been much work in evaluating if they can generalize to new instructions (Weller et al., 2020; Wang et al., 2022c; Ouyang et al., 2022), if we can train them to follow instructions without human-annotated data (Wang et al., 2022b; Qin et al., 2023), and applying them to various domains (Zhao et al., 2021; Singhal et al., 2023; Shaghaghian et al., 2020). As the IR community uses LLMs in their models, we seek to broaden the scope of IR to include instructions, aligning it with the broader NLP community.
2.3 Instructions for Retrieval
Using instructions in retrieval is a nascent area of exploration. Su et al. (2022) and Asai et al. (2022) were two of the earliest works that trained a retrieval model to use instructions along with the query. However, these instructions are typically very short, such as “Retrieve a Wikipedia paragraph that answers this question." Recent work incorporates instructions in smaller models (Xiao et al., 2023; Chen et al., 2023, 2024) as well as others which use Mistral (Jiang et al., 2023) as the backbone of a larger retrieval model that can use instructions: GritLM (Muennighoff et al., 2024) trains Mistral to do both generation and embedding, while Wang et al. (2023) uses Mistral for embeddings only.
Despite this flurry of activity, these efforts do not have an explicit instruction-related retrieval benchmark to evaluate on. Instead, they evaluate on standard retrieval benchmark suites such as MTEB (Muennighoff et al., 2022) and BEIR (Thakur et al., 2021) which do not contain instructions. Thus, these newer instruction-IR models hand-write a few instructions, where typically each instruction is applied to an entire dataset, irrespective of the query. This makes these instructions generic: focused only on the task format, format of the “document" (paragraph, sentence, etc.), and the broad domain. Note that because of this, no current instructions contain any extra background information or negation (Weller et al., 2024) which are commonly found in real-world instructions (Table 1).
In work concurrent to ours, Oh et al. (2024) also propose a dataset to evaluate instructions in retrieval models. Their dataset uses the MSMarco collection, and differs in several crucial aspects: it only has one relevant document per query (e.g. sparsely judged), is GPT-4 generated and validated, focuses on the background of the user (“I am a school teaching looking for …"), and evaluates using the lowest score over $N$ instructions for the same query (measuring robustness). In contrast, we use highly-judged corpora to ensure we can measure recall, use real-world human generated instructions, have human-validated relevance judgements, propose a new paired evaluation protocol, and provide a training dataset and model for teaching instruction-following.
Dataset # $Q$ $|Q|$ $|I|$ Rel. D/Q # $Q$ $|I|$ Rel. D/Q
TREC News ’21 (Soboroff et al., 2020) 50 15.3 40.1 50.1 32 46.9 19.2
TREC Core ’17 (Allan et al., 2017) 50 16.6 44.0 180.0 20 53.5 32.7
TREC Robust ’04 (Voorhees, 2005) 249 11.9 68.2 69.9 52 75.2 19.8
Table 2: FollowIR evaluation set statistics before (left) and after (right) annotation. We use a subset of the queries in three popular TREC tracks for variety in queries and documents. $|Q|$ is the word length of the queries and $|I|$ is the word length of the instructions. Rel. D/Q indicates the number of relevant annotated documents in the collection, excluding irrelevant annotations. As designed, there are less relevantly-judged documents in the FollowIR portion (as the annotations change the relevance of documents on purpose for evaluation).
3 Building FollowIR
We build FollowIR by starting with three existing TREC collections: TREC News 2021 that uses the Washington Post v4 corpus (Soboroff et al., 2020), TREC Robust 2004 using the Disks 4 and 5 collections of various news articles (Voorhees, 2005), and TREC Common Core 2017 using the New York Times Annotated corpus (Allan et al., 2017). Each of these was human annotated to include hundreds of annotations per query (see Table 2), with around 50-180 relevant documents per query on average (and many more not-relevant annotations).
Each of these TREC tracks includes instructions for the human annotators that we use as the instructions. Although using these alone can provide some indication of how well models can follow instructions, it doesn’t explicitly test their instruction following ability. To more carefully isolate this ability in our benchmark, we test whether models can respond to small changes in the instruction.
To accomplish this, we ask two expert human annotators to modify the TREC instructions. However, doing this in a naive way would require re-annotating all the document judgements, a non-trivial task requiring immense annotation efforts. [^3] NIST spends roughly 1-2 million per year: https://trec.nist.gov/pubs/2010.economic.impact.pdf We crucially ask the annotators to make the instructions more specific by including additional aspects to the relevance definition or to include relevance instructions with negation. This decreases the set of documents to re-annotate to only those documents deemed relevant by the original TREC annotators (as it is then a subset of the original relevant documents by definition). This makes the annotation tractable, with only dozens or hundreds of documents to re-annotate per query instead of needed to re-annotate the whole collection (e.g. thousands).
We annotate a subset of the original TREC queries due to cost and overlap: we sample 50 queries from TREC Robust 2004 that do not overlap with TREC Common Core (as Common Core used 50 queries from Robust04 on a new collection), and 30 queries from TREC News 2021. Table 2 shows dataset statistics of judged documents and the final benchmark size. Annotators were asked to change the instructions so that the number of relevant documents was cut in half, thus including a sizeable number of changed relevance judgements. We note that although the number of queries seems small by NLP standards, 30-50 queries is both effective (Webber et al., 2008) and standard in the IR community due to the expense of careful annotation over many documents per query.
Due to differences in retriever quality, if we evaluate by searching over the full collection, each model will retrieve a different number of relevant documents. However, because we evaluate instruction following based on changing the document relevance, models that do poorly in the initial retrieval will have fewer documents which change relevance in the instruction-following evaluation. To rectify this, we instead turn to a reranking task where we include all relevant documents, and use a pool of five models [^4] We use BM25, BGE-base, E5-base-v2, TART-Contriever, and INSTRUCTOR-XL to select the top non-relevant documents. To be able to freely distribute the data due to fair-use laws, we chunk the documents into 400-word passages with 200-word overlap and select the highest scoring passages using MaxP (Dai & Callan, 2019). This enables us to distribute our data, which we do by extending the MTEB evaluation framework (Muennighoff et al., 2022).
3.1 Evaluation Metrics for FollowIR
Our benchmark provides two ways of measuring instruction following: (1) standard retrieval metrics when using the instructions with the queries and (2) pairwise evaluation of instruction following. (1) We use typical IR evaluation metrics but use the instruction along with the query: these metrics are mean average precision (MAP) for Core17/Robust04 and normalized discounted cumulative gain at 5 (nDCG@5) for News21. For (2) we use our novel pairwise evaluation metric that measures the delta in scores when following the modified instructions instead of the original. [^5] Note that we do not show standard retrieval results on the modified instruction’s relevant document set, as standard retrieval scores cannot be directly compared across different query relevance annotations (qrels).
Our new pairwise evaluation metric, $p$-MRR, measures rank-wise changes between queries. In developing this metric we had the following desiderata: it should compare the results of the original instruction to those of the new instruction, it should have a standardized range from worst possible change in instruction-following score (i.e. -1) to best possible instruction-following score (i.e. 1) with an option for no change when using different instructions (i.e. 0), and finally should take into account the document rank so that changes from rank 1 to rank 2 are more prominent than changes from rank 99 to 100. Given the above qualifications, we use the following equation applied to each relevant document per query (where MRR is mean reciprocal rank, $R_{og}$ is the rank of the document when using the original instruction and $R_{new}$ is the new rank):
[…]
For the final score, we average first within a given query and then over all queries in the corpora – i.e. macro-averaging across queries, to handle the different number of relevant documents per query.
4 Evaluating Instruction Following
In this section we describe the models we evaluate, results of those models on FollowIR, and ablations performed to better understand the behavior of current models.
4.1 Evaluation Settings
We evaluate a wide variety of IR models (trained with and without instructions), including neural models ranging from 100 million to 7 billion parameters. We evaluate on the original TREC instructions in the FollowIR benchmark and then on the new instructions, showing both standard IR metrics and the new pairwise metric $p$-MRR. We group models into four categories:
No Instructions in Training
These retrieval models did not see instructions in training and typically aren’t given them: this includes Contriever (Izacard et al., 2021), E5 (Wang et al., 2022a), MonoBERT Nogueira et al. (2019), MonoT5 (Nogueira et al., 2020), and BM25 (Robertson et al., 1995).
Instructions in IR Training
Most retrieval models using instructions received roughly one instruction per retrieval dataset, which generally defined the domain (e.g. “Financial"), document size (sentence, passage, etc.), and task format. This includes INSTRUCTOR models (Su et al., 2022), the bi-encoder TART model trained from Contriever (Asai et al., 2022), the reranker TART trained from FLAN-T5 (Chung et al., 2022), E5 Mistral-Instruct (Wang et al., 2023), and GritLM (Muennighoff et al., 2024). We also include BGE models (Xiao et al., 2023) in this category, although they are trained with only one instruction total for each broad task (retrieval, clustering, etc.).
API Models
We use two of the best performing API embedding models: Cohere’s v3 English and OpenAI’s Text-Embedding-v3-Large. It is unknown what these model’s training procedures were – including if they were trained on instructions or not – thus we place them in a distinct category.
Instruction-Tuned LLMs
We also evaluate several instruction-tuned LLMs to be used as rerankers, including FLAN-T5 (Chung et al., 2022), Llama v2 (Touvron et al., 2023b), and Mistral-Instruct-v0.2 (Jiang et al., 2023). We evaluate these models in the same fashion as MonoT5 rerankers, comparing the true and false tokens. Note that these models were not trained on any retrieval-specific data.
Robust04 News21 Core17 Average
Model mAP $p$-MRR nDCG $p$-MRR mAP $p$-MRR Score $p$-MRR
No-Instruction IR
BM25 12.2 -2.8 21.3 +2.5 8.1 -0.7 13.9 -0.3
E5-base-v2 14.5 -6.8 21.6 -4.1 14.0 -2.9 16.7 -4.6
E5-large-v2 18.1 -4.1 24.9 -2.2 17.0 +0.1 20.0 -2.1
Contriever 20.3 -6.1 24.0 -1.8 15.3 -2.5 19.9 -3.5
MonoBERT 21.5 -9.7 26.3 -4.4 18.4 -1.3 22.1 -5.1
MonoT5-base 16.3 -5.8 11.9 -1.2 12.2 -3.5 13.5 -3.5
MonoT5-3B 27.8 +5.6 18.6 +7.5 18.1 +1.7 21.5 +4.9
Instruction-IR
BGE-base 17.5 -6.4 23.8 -0.2 14.6 -2.7 18.6 -3.1
BGE-large 18.1 -7.8 26.4 +0.1 15.0 +0.1 19.8 -2.5
TART-Contriever 14.1 -7.8 21.9 +0.0 12.4 -1.3 16.1 -3.0
INSTRUCTOR-base 14.4 -5.6 16.3 -2.5 14.7 -2.2 15.1 -3.4
INSTRUCTOR-xl 15.5 -2.1 14.6 -4.3 14.4 -0.6 14.8 -2.3
TART-FLAN-T5-xl 25.2 -0.8 20.3 -1.1 17.0 +2.8 20.8 +0.3
GritLM-7B 29.0 -1.4 25.2 +2.1 20.8 +2.6 25.0 +1.1
APIs
Cohere v3 English 22.9 -3.3 23.6 -3.1 20.6 +2.7 22.4 -1.2
OpenAI v3 Large 27.9 -5.7 30.0 -3.3 21.4 -0.2 26.4 -3.1
Instruct LMs
FLAN-T5-base 6.8 +5.0 2.2 +1.1 6.5 -3.2 5.2 +1.0
FLAN-T5-large 15.1 +4.0 8.5 +7.7 11.5 +1.2 11.7 +4.3
Llama-2-7B-chat 6.9 +1.6 13.3 +2.1 5.4 +3.6 8.5 +2.4
Mistral-7B-instruct 24.1 +12.2 22.9 +10.5 19.6 +13.4 22.2 +12.0
FollowIR-7B 25.9 +13.6 25.7 +10.8 20.0 +16.3 23.9 +13.6
Table 3: Evaluating instruction-following on FollowIR. $p$-MRR is a new pairwise evaluation metric measuring instruction following when instructions change, ranging from -100 to 100 (higher is better). We see that the only models that show any success at following instructions are large models (3B+ parameters) or instruction-tuned LLMs that haven’t been trained on retrieval tasks.
4.2 Results
Table 3 shows the main results on FollowIR, with the standard IR score shown (either MAP or nDCG@5) as well as the pairwise evaluation metric, $p$-MRR.
No-Instruction IR Models
We see that the no-instruction models range in standard IR metrics (in terms of nDCG@5 and MAP) but generally have negative scores for $p$-MRR (up to -9.7). The only non-instruction model to score positively on average is MonoT5-3B.
Instruction IR Models
We again see that these models have generally negative scores, with the exception being TART-FLAN-T5-xl (with scores averaging roughly zero) and GritLM which has slightly positive scores for two of the three datasets.
Figure 2: Score difference when changing from no instructions to using instructions formatted as keywords, short text, or the full text. We see that models that can correctly use instructions see gains with the additional information while most other models see decreasing performance as instruction length increases. This indicates that most models struggle with instruction length, using these instructions for keyword search rather than as instructions that define document relevance.
API Models
We see that the API models perform strongly in terms of standard IR metrics, with OpenAI’s model performing the highest overall. However, both OpenAI and Cohere’s models perform poorly at instruction-following with negative scores (-3.1 and -2.8 on average, respectively).
Instruct-Tuned LLMs
In contrast to the previous results, all instruction-tuned LLMs show positive results for instruction following, although they have the widest range of performance using standard IR metrics (ranging from very poor scores to some of the higher scores). We see that the best performing model in this category is FollowIR–7B, which we describe in more detail in Section 5.
Overall
We see that the only models that show positive results at following instructions are either larger IR models (3B+ parameters) or those that have been explicitly trained to follow instructions (FLAN-T5) without retrieval-specific training data. This aligns with work in the natural language processing community which has shown that the instruction-following ability improves with scale (Brown et al., 2020) and instruct-tuning (Longpre et al., 2023).
4.3 Analysis
Why do so many models fail to correctly follow instructions when they do well on typical IR metrics such as nDCG and MAP? We answer this question by ablating several components that may impact results: (1) whether IR models are not used to instructions that cannot be used for keyword search and (2) whether they are unused to the length of the longer instruction (as instruction-IR models have been trained on much shorter instructions).
We compare the original query-only result to those where we additionally give the model either the full instruction, a shorter instruction, or keywords from the instruction. We gather these short instructions and keywords by prompting GPT-4-Turbo-1106 to generate them from the original full instruction (for TREC data) or the original short instruction used by the model (for BEIR data). For the full prompt text, please see Appendix E.
We show results for these ablations in Table 2, where positive scores indicate that adding information improves the model while negative scores indicate a drop in performance. We see a consistent trend where models that did poorly on longer instructions perform better on keywords and shorter instructions than with the full instruction. However, models that are able to follow instructions see better results with the additional information, on average.
These results show that models are (1) using the instruction text as keywords and (2) are unused to longer (with less relevant) text that is in the instructions.
We also confirm that these results hold on datasets outside of TREC collections and show results on three BEIR datasets: SciFact, NFCorpus, and FiQA. We show in Table 4 the original score (using the short instructions from their papers) and the change in score when using just keywords from the instruction (again extracted from GPT-4). We show results only for models which performed poorly for instruction-following. We see that the scores for keywords vs the short instruction are generally similar, with most models seeing a change of around $\pm$ 1 point, except for the strongest instruction-following model, E5-Mistral, seeing a larger drop on some datasets.
Overall
We find overall (on both TREC and BEIR datasets) that models use instructions for keyword matching and are unused to longer instructions that may contain slightly less relevant words.
SciFact NFCorpus FiQA
Model OG $\Delta$ w/Key. OG $\Delta$ w/Key. OG $\Delta$ w/Key.
No-Instruction
BM25 67.9 -1.7 32.2 -5.1 23.6 -1.6
E5-base-v2 71.9 -2.7 35.4 -2.5 39.9 -0.4
Contriever 64.9 +0.4 31.7 +0.0 24.5 -3.2
MonoT5-base 73.1 -0.6 35.6 -0.9 41.2 -0.3
Uses Instruction
TART-Contriever 67.6 -0.3 33.4 -5.3 31.8 -0.4
INSTRUCTOR-base 57.8 +1.0 31.6 -0.4 39.2 -0.1
BGE-base 73.2 -0.5 35.5 +0.0 40.8 -2.3
TART-FLAN-xl 74.2 +1.6 33.9 +0.4 39.6 -0.3
INSTRUCTOR-xl 62.4 +0.2 36.0 -0.6 46.9 +0.8
E5-Mistral 77.1 -5.1 38.8 +0.3 56.7 -6.5
Table 4: Ablation on BEIR benchmarks for models that do poorly with longer instructions, comparing their original short instructions vs domain keywords extracted from those instructions (see Appendix D for a list). If models had learned to use the instructions correctly we would see a divergence between the behavior of instruct and non-instruct models, however, for both we see that using keywords instead of the instruction results in comparable performance ($\pm$ one point).
5 Teaching Instruction Following
Is it possible to improve model performance in following instructions? We seek to show that fine-tuning on a training set of longer instructions can provide a method for doing so.
We start by gathering a training set to teach models. We collect all TREC narratives (i.e. instructions) from tasks not in FollowIR, consisting of 1836 pairs of queries and narratives. However, we note that this does not provide any positive or negative documents for fine-tuning.
In order to obtain documents for training, we prompt GPT-3.5-Turbo-1106 to generate relevant and not-relevant documents, generating roughly two relevant and non-relevant instances per query. However, these synthetic documents are noisy and contains errors w.r.t. the labels – to remedy this, we perform a round of filtering and use the best performing open-source model from Table 3 (Mistral-7B-Instruct-v0.2) to score each of the generated documents according to the instruction. We then filter the documents according to whether Mistral correctly predicts the generated label, and finally balance the relevant and non-relevant samples, choosing only one relevant and non-relevant document per query. Our total is $\sim$1800 training instances on $\sim$1200 unique query/instructions pairs.
We then train our instruction-following model, FollowIR-7B, by fine-tuning Mistral-7B-Instruct-v0.2 on our data using the Llama-Factory framework (Hiyouga, 2023) with LoRA (Hu et al., 2021). Full training hyperparameter details are found in Appendix A.
When we evaluate this model on FollowIR (Table 3), we find that the scores consistently improve. Compared to the original Mistral-7B-Instruct-v0.2, our model improves on both standard IR metrics (+7.6% relative improvement) and on instruction following (+13.3% relative improvement). Thus, we can see that it is possible to train IR models to be better instruction followers.
6 Conclusion
Despite the use of LLMs as the backbone of neural retrieval models, most existing IR models do not take instructions that define document relevance. Further, there is no existing resource that measures how well retrieval models can follow instructions. We build a new benchmark that explicitly measures the instruction following ability of retrieval models and find that nearly all retrieval models do not follow instructions, with the exception of very large models (3B+ parameters) or instruction-tuned LLMs that typically are not used for retrieval. However, we show that it is possible to improve their instruction following ability, and build and release a training corpus for teaching retrieval models to follow instructions. Our new model, FollowIR-7B, shows improvement on both standard retrieval metrics as well as in instruction following. We hope that these resources will allow the community to develop more capable instruction-following retrieval models that can quickly adapt to a relevance definition given flexible natural language text.
7 Limitations
Reranking vs Full Retrieval
As evaluating instruction following requires evaluating the documents which changed relevance, we cannot use the full collection for retrieval (as each retriever finds different relevant documents by design). Further, due to licensing restrictions, we cannot distribute the full corpora from the TREC tracks – thus we distribute passages due to fair use laws. However, we show full corpus retrieval results for a subset of models in Appendix C and note similar trends.
Possible Errors
Our work is built on the TREC document collections and judgements, as well as new annotation efforts. We do not check for potential errors in the TREC annotations, and our newly gathering annotations may have small errors. Despite these caveats, we see that our dataset still provides a useful evaluation setup for measuring instruction following.
Acknowledgments
Orion Weller is supported by a National Science Foundation Graduate Research Fellowship Program grant. We also thank Marc Marone for their feedback and comments that helped improve this work.
Appendix A Hyperparameters for Fine-Tuning Mistral
We fine-tune Mistral-7B-Instruct-v0.2 using the Mistral template for 8 epochs. We use LoRA and tune the “q_proj,v_proj,o_proj,k_proj" parameters, using a rank of 8 and an alpha of 16. We use bfloat16, a batch size of 32, and a learning rate of 3e-5.
Appendix B Hyperparameters for Inference
We use default parameters for inference, taken from the original code of the authors of the papers we use (from their MTEB evaluations).
Appendix C Full Retrieval Results
In Table 5 we show results for models searching on the full collections of the TREC tasks included in FollowIR. Note that because each model retrieves different relevant documents, the instruction-following evaluation has a different set of instances that each model is evaluated on (as it can only be evaluated on documents it retrieved that then become not-relevant).
Robust04 News21 Core17
(mAP) (nDCG@5) (mAP)
Model OG $\Delta$ OG $\Delta$ OG $\Delta$
No
Instruct
BM25 21.4 -1.2 30.1 +5.3 16.8 -0.2
E5-base-v2 22.7 -7.0 33.6 +1.8 19.7 -3.0
Contriever 19.2 -7.7 22.5 +9.0 22.6 -7.6
Uses
Instruct
TART-Contriever 25.5 -10.1 40.0 -5.0 22.6 -7.6
BGE-base 23.6 -3.1 36.5 -7.8 23.0 -2.1
INSTRUCTOR-base 22.5 -2.2 33.3 -2.8 20.0 -0.2
INSTRUCTOR-XL 30.4 -3.1 38.1 -0.1 29.9 -2.8
Table 5: FollowIR scores on the full retrieval collection (thus rerankers are not included). As the base score is different, there are different numbers of relevant documents they are being evaluated on for $p$-MRR. Thus, we only report the original (no-instruction) score and the delta when using the TREC instructions. We note that it shows similar results to the main text – retrieval models are not effectively using instructions and see performance degradations with longer text.
Appendix D Keywords used for BEIR experiments
GPT-4-Turbo-1106 extracted the following keywords (Table 6) from the instructions these models used, which generated the results in Table 4.
Model Dataset Keywords
BM25/Contriever/E5/MonoT5 FiQA Finance web
BM25/Contriever/E5/MonoT5 SciFact science paper verify
BM25/Contriever/E5/MonoT5 NFCorpus medicine relevant
TART-dual FiQA financial web
TART-dual SciFact scientific paper verify
TART-dual NFCorpus scientific paper paragraph
INSTRUCTOR-base FiQA financial supporting:
INSTRUCTOR-base SciFact scientific supporting passage:
INSTRUCTOR-base NFCorpus medicine relevant
BGE-base FiQA relevant passages:
BGE-base SciFact relevant passages:
BGE-base NFCorpus relevant passages:
INSTRUCTOR-xl FiQA finance supporting:
INSTRUCTOR-xl SciFact scientific supporting passages:
INSTRUCTOR-xl NFCorpus nutrition facts public medical:
E5-Mistral FiQA financial replies
E5-Mistral SciFact scientific
E5-Mistral NFCorpus retrieve relevant
TART-T5-FLAN-xl FiQA financial web
TART-T5-FLAN-xl SciFact scientific paper verify
TART-T5-FLAN-xl NFCorpus Scientific paper paragraph
Table 6: Keywords used for the BEIR keyword analysis. Note that non-instruction models received the keywords used in INSTRUCTOR-base and TART-dual (as shown in the table).
Appendix E Prompts Used
We use these prompts for generating the short instructions, the keywords, and the synthetic documents. The examples used in the prompt for the “Full Instructions to Short Instructions" prompt were partially created by the authors, as only the short instructions were provided by TART/INSTRUCTOR.
I have instructions that are specific to a question, but I need your help abstracting them to a general task format that I can give to someone else. I need you to turn them into an abstract command that just describe the general abstract task instead (e.g., where the data is from, what the type of document looks like). It is crucial that you read and follow these instructions, you will get a large bonus if you are successful (\$200).
The abstract command should only mention the **task format**. Do **not** refer to any entities or specific text in the original instruction. Your response should be around 10 words. The command should be as if you were speaking to another human.
## Examples
### Example 1:
Original Instruction: A relevant document would provide information about the whole blood-base perfusate and whether or not it provides superior preservation of myocardial function during ex vivo heart perfusion. This may include research experiments, commentary, or survey/review papers. Information about whole blood-base perfusate alone is not relevant, unless it also mentions it’s effect on myocardial function during ex vivo heart perfusion.
Abstract Command: Help me to find a highly related PubMed paper to answer this question.
### Example 2:
Original Instruction: A relevant document will contain information that about the right of way in international waters that can be used to determine who should have the right of way in a given situation. For example, it should contain instances about who is at fault in an accident, if it depends on the size of the boat, or details about how this differs according to nationality. Especially relevant are documents describing who is at fault in a crash situation.
Abstract Command: Retrieve a Wikipedia paragraph that answers this question.
### Example 3:
Original Instruction: A relevant instance will be a question that is semantically equivalent to the query given. For example, it may contain different lexical words or be a paraphrase of the other, but the underlying meaning will be the same. If the instance is not semantically the same as the query, it is irrelevant.
Abstract Command: Check if a Quora question is duplicated with this question.
### Example 4:
Original Instruction: A relevant document would include details about the timing of medicare and what age patients can start using it for healthcare. It may include information about laws, insurance, or other details that describe the age the medicare begins. Less relevant are documents talking about potential laws or facts about medicare that do not answer the question of what age medicare begins. Just the mention of an age and when to start medicare would be relevant.
Abstract Command: I want to know the answer to the question. Can you find good evidence on the web?
Now you can easily see that the abstract command is vague and describes only a short command about how to get the information you need. Follow this exactly – do not reference specifics (like in the above, "international waters" and "medicare" are not included in the abstract command). You should instead keep the abstract command vague and well, abstract about the task only. Use the word "question".
## Your turn
Original Instruction: FILL_TEXT_HERE
Abstract Command (remember to use the word "question"):
I have instructions that are specific to a style of retrieval, but I want you to instead just focus on the relevant keywords that are in these instructions. Your job is to return a list of these keywords that are relevant in the query. There are probably one or two relevant keywords to extract only. ## Examples ### Example 1: Instruction: Help me to find a highly related PubMed paper to answer this question. Keywords: ["PubMed"]
### Example 2: Instruction: I want to find an answer for this Trivia question. Can you find some paragraphs that provide evidence from Wikipedia? Keywords: ["Trivia", "Wikipedia"]
### Example 3: Instruction: Check if a Quora question is duplicated with this question. Keywords: ["Quora", "duplicated"]
### Example 4: Instruction: I want to find a related question asked in StackExchange. Can you find one for me? Keywords: ["related", "StackExchange"]
## Your turn Instruction: FILL_TEXT_HERE Keywords (either one or two keywords, that are not "documents", "questions", "answer", or "articles"):
I need you to annotate some data for my business and it is super important that you follow instructions precisely or you will be fired. Given a Google search query and instructions regarding what makes a document relevant, I need you to write two documents: one that would be relevant and one that would not. Search: TITLE_HERE Instructions: NARRATIVE_HERE I need some different options to choose from, so give me three **different** options for both a relevant document and an irrelevant document. They should be **long** paragraph-sized documents ($\sim$300 words each), one on each line. If there is no negation in the instructions, your irrelevant document should be slightly off topic: